Zeronknight (zeronknight):

Simplify the following ratio. 12 to 4 A. 1 to 3 B. 1 to 4 C. 3 to 1

Zeronknight (zeronknight):

How do I simplify ratios?...

OpenStudy (anonymous):

you divide both numbers by the same number. Do you have any guesses?

Zeronknight (zeronknight):

12 / 4 = 3 4 / 4 = 3

Zeronknight (zeronknight):

So C

Zeronknight (zeronknight):

@emandile57

Zeronknight (zeronknight):

NOT 3. i meant 1

Zeronknight (zeronknight):

4 / 4 = 1

OpenStudy (anonymous):

C is correct :)

Zeronknight (zeronknight):

Alright thanks! :)

OpenStudy (anonymous):

no problem